Cost Information and FIFO Gunnison Company had the following equivalent units schedule and cost information for its Sewing Department for the month of December: Direct Materials Conversion Costs Units started and completed 40,000 40,000 Add: Units in beginning work in process × Percentage complete: 5,000 × 0% direct materials — 5,000 × 50% conversion Costs 2,500 Add: Units in ending work in process × Percentage complete: 12,000 × 100% direct materials 12,000 — 12,000 × 35% conversion Costs — 4,200 Equivalent units of output 52,000 46,700 Costs: Work in process, December 1: Direct Material $65,000 Conversion Costs 10,000 Total work in process $75,000 Current costs: Direct Material $728,000 Conversion Costs 140,100 Total current costs $868,100 Required: 1. Calculate the unit cost for December, using the FIFO method. $ per equivalent unit 2. Calculate the cost of goods transferred out, calculate the cost of EWIP, and reconcile the costs assigned with the costs to account for. Cost of goods transferred out $ Cost of EWIP $ Cost to account for: BWIP $ Current (December) Total $ 3. What if you were asked for the unit cost from the month of November? Calculate November's unit cost. $ per equivalent unit
Answer #1

solution Materials = 728000 / 52000 = 14 Conversion = 140100 / 46700 = 3 Cost per equivalent unit = 17 Cost of goods transfer